# Packing Input Frame Context in Next-Frame Prediction Models for Video Generation

### v2

@@ 3.3 Anti-drifting Sampling
Figure 2: Visualization of Sampling Approaches. We present sampling approaches to generate frames in different temporal orders. The shadowed squares are the generated frames in each iteration, whereas the white squares are the iteration inputs. Note that only the vanilla approach relies on causal inputs, and both the anti-drifting and inverted anti-drifting approaches receive bi-directional inputs.
Drifting is a common problem in next-frame prediction models where visual quality degrades as video length increases. While the underlying cause remains an open research problem, we observe that drifting only happens in causal sampling (i.e., when models only access past frames). We show that providing access to future frames (even a single future frame) will get rid of drifting. We point out that bi-directional context, rather than strictly causal dependencies, might be fundamental for maintaining video quality.
@@
The vanilla sampling method shown in Fig. 2-(a), i.e., iteratively predicting future frames, can be modified into Fig. 2-(b), where the first iteration simultaneously generates both beginning and ending sections, while subsequent iterations fill the gaps between these anchors. This bi-directional approach prevents drifting since the ending frames are established in the first iteration, and all future generations attempt to approximate them.
@@
We discuss an important variant by inverting the sampling order in Fig. 2-(b) into Fig. 2-(c). This approach is effective for image-to-video generation because it can treat the user input as a high-quality first frame, and continuously refines generations to approximate the user frame (which is unlike Fig. 2-(b) that does not approximate the first frame), leading to overall high-quality videos.
All three methods in Fig. 2-(a,b,c) can generate videos of arbitrary length. Method (a) achieves this through direct iterative generation, while in methods (b) and (c), we can dynamically move the ending sections (or generated frames) to greater distances as our generated frames approach them. Alternatively, in practice, setting a sufficiently large time range (e.g., 1 minute) in the first iteration typically satisfies practical requirements.
@@
These sampling methods require modifications to RoPE to support non-consecutive phases (time indices of frames). This is achieved by skipping the non-queried phases (indices) in the time dimension.
@@ 4.1 Ablative Naming
Kernel (simplified): For simplicity, since kernels that are multiples of $(1,2,2)$ are commonly used, we use abbreviated notation such as k1 that only denotes the temporal dimension. Specifically, k1 represents k1h2w2 (the kernel $(1,2,2)$), k2 represents k2h4w4 (the kernel $(2,4,4)$), k4 represents k4h8w8 (the kernel $(4,8,8)$), etc.
@@
Packing: The notation f16k4f4k2f1k1 shows a way to encode $19$ contiguous frames with three compression levels: the first $16$ frames are encoded by the kernel k4 (the kernel $(4,8,8)$), the next $4$ frames are encoded by the kernel k2 (the kernel $(2,4,4)$), and the last $1$ frame is encoded by the kernel k1 (the kernel $(1,2,2)$).
@@
With the above naming convention, we can represent all ablative structures in a compact form. Note that this naming also implies the sampling approach as discussed in Section 3.3. Consider the three similar structures and their sampling:
@@ 4.2 Base Model and Implementation Details
We implement FramePack with Wan and HunyuanVideo. We implement both the text-to-video and image-to-video structures, though both are naturally supported by next-frame-section prediction models and do not need architecture modifications.
The Wan base model is the official Wan2.1 model. The HunyuanVideo model is an improved version of official HunyuanVideo models to match the capability of Wan2.1, with the modifications: (1) adding the SigLip-Vision model google/siglip-so400m-patch14-384 as a vision encoder, (2) removing the reliance on Tencent’s internal MLLM, (3) freezing LLama3.1 as a pure text model and ignoring its multi-modality, and (4) continued training on high-quality data.
Note that the numerical results presented in this paper should not be interpreted as direct comparisons between Wan and HunyuanVideo, as both were trained with similar computational budgets while Wan’s larger model size means fewer relative resources (e.g., batch size). Both models demonstrate comparable quality after sufficient training. We recommend HunyuanVideo as the default configuration for more efficient training and faster inference.
Dataset
We follow the guidelines of LTXVideo [13]’s dataset collection pipeline to gather data at multiple resolutions and quality levels. All data are filtered using quality measurements and motion scores to ensure a high-quality diverse distribution. We use aspect ratio bucketing for multi-resolution training with a minimum unit size of $32$ pixels. For example, the buckets at 480p resolution include: $(416,960)$, $(448,864)$, $(480,832)$, $(512,768)$, $(544,704)$, $(576,672)$, $(608,640)$, $(640,608)$, $(672,576)$, $(704,544)$, $(768,512)$, $(832,480)$, $(864,448)$, and $(960,416)$.
Small-scale Training
FramePack achieves a batch size of 64 on a single 8×A100-80G node with the 13B HunyuanVideo model at 480p resolution (without using any image-based workaround training that many community LoRAs are built on top of). This batch size is comparable even to image diffusion models like the 12B Flux, making FramePack suitable for personal or laboratory-scale training and experimentation.
Training
We conduct all experiments using H100 GPU clusters. Each FramePack variant in the ablation studies consumes approximately 48 hours of training, while the final models are trained for one week. We use the Adafactor optimizer with a fixed learning rate of 1e-5 without learning rate scheduling. The gradient norm clip is set to 0.5.
Lower flow shift
Since the next-frame-section prediction methods generate smaller 3D tensors at each inference compared to full-video diffusion, the models can be trained with much lower flow shift values. We use Flux’s dynamic flow shift to train all models. We find that this leads to sharp and clean results that are closer to real videos.
4.3 Evaluation Metrics
@@
4.3.1 Multi-dimension Metrics
Multiple metrics for video evaluations are consistent with common benchmarks, e.g., VBench [18], VBench2 [61], etc. All scores are normalized using the same numerical system with VBench.
Clarity: The MUSIQ [22] image quality predictor trained on the SPAQ [10] dataset. This metric measures artifacts such as noise and blurring in the generated frames.
Aesthetic: The LAION aesthetic predictor [33]. This metric measures the aesthetic values perceived with a CLIP-based estimator.
Motion: The video frame interpolation model [17] modified by VBench to measure the smoothness of motion.
Dynamic: The RAFT [37] modified by VBench to estimate the degree of dynamics. Note that the “dynamic” metric and “motion” metric represent a trade-off, e.g., a still image may rank high on motion smoothness but will be penalized by low dynamic degrees.
Semantic: The video-text score computed by ViCLIP [42]. This metric measures the overall semantic consistency between the generated video and the input text.
Anatomy: The ViT [9] pretrained by VBench for identifying the per-frame presence of hands, faces, bodies, etc.
Identity: The facial feature similarity using ArcFace [7] with face detection by RetinaFace [8] to measure the identity consistency.
Table 1: Ablation study. We evaluate different FramePack configurations across multiple global metrics, drifting metrics, and human assessments. The table is divided into three groups based on sampling approach: vanilla sampling (top), anti-drifting sampling (middle), and inverted anti-drifting sampling (bottom). The tests are conducted with HunyuanVideo as base. Bests in bold. ELO differences within ±$\pm 16$ are considered ties. Variant Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ td_f8k8f4k4f2k2f1k1_g9 67.33% 65.94% 94.53% 92.91% 20.06% 66.97% 71.72% 3.25% 3.45% 7.45% 16.56% 1084 8 td_f64k8f16k4f4k2f1k1_g9 67.71% 66.71% 95.09% 93.91% 21.39% 66.56% 71.50% 3.22% 3.38% 7.25% 16.43% 1068 9 td_f512k8f64k4f8k2f1k1_g9 67.74% 66.44% 94.27% 92.91% 21.90% 67.73% 71.35% 3.18% 3.32% 7.05% 15.95% 1088 8 td_f512k16f64k8f8k2f1k1_g9 66.13% 64.52% 94.77% 94.18% 19.21% 65.72% 71.51% 3.25% 3.45% 7.85% 17.62% 1072 9 td_f16k4f4k2f1k1_g9 67.37% 65.05% 95.97% 91.66% 19.15% 65.38% 71.73% 3.22% 3.48% 6.65% 17.36% 1076 9 td_f16k4f2k2f1k1_g1 65.81% 64.39% 94.91% 94.92% 19.20% 64.16% 69.70% 3.43% 3.77% 9.81% 20.89% 1036 11 td_f16k4f2k2f1k1_g4 66.57% 64.99% 94.00% 82.85% 19.40% 65.97% 69.06% 3.36% 3.68% 8.55% 19.09% 1052 10 td_f16k4f2k2f1k1_g9 67.15% 65.29% 94.08% 92.97% 20.73% 66.46% 71.08% 3.18% 3.42% 7.45% 18.05% 1080 9 tc_f16k4f2k2f1k1_g9 67.62% 65.07% 94.02% 90.33% 21.71% 71.70% 71.01% 3.15% 3.25% 7.25% 17.21% 1092 8 ta_f16k4f2k2f1k1_g9 67.02% 66.44% 94.11% 91.09% 21.48% 71.92% 72.18% 3.12% 3.18% 7.05% 17.66% 1096 8 td_f8k8f4k4f2k2f1k1_g9_x_f1k1 68.46% 66.95% 96.46% 75.55% 22.88% 85.10% 75.84% 2.95% 2.85% 5.95% 15.87% 1132 4 td_f64k8f16k4f4k2f1k1_g9_x_f1k1 68.21% 67.75% 95.74% 85.97% 22.79% 81.09% 76.29% 2.92% 2.98% 5.95% 15.99% 1136 4 td_f512k8f64k4f8k2f1k1_g9_x_f1k1 68.62% 67.72% 95.05% 76.50% 22.44% 82.01% 76.65% 2.88% 2.92% 5.75% 15.94% 1116 5 td_f512k16f64k8f8k2f1k1_g9_x_f1k1 68.35% 67.89% 97.73% 76.48% 22.75% 78.40% 76.16% 2.85% 2.85% 5.55% 14.04% 1120 5 td_f16k4f4k2f1k1_g9_x_f1k1 68.42% 67.59% 96.74% 74.37% 23.69% 79.14% 77.37% 2.82% 2.78% 5.35% 14.90% 1140 4 td_f16k4f2k2f1k1_g1_x_f1k1 65.32% 67.97% 95.26% 81.69% 19.97% 74.57% 77.93% 2.78% 2.72% 4.95% 14.80% 1084 7 td_f16k4f2k2f1k1_g4_x_f1k1 69.92% 67.49% 97.84% 71.90% 21.12% 74.84% 77.53% 2.75% 2.65% 4.95% 14.98% 1100 6 td_f16k4f2k2f1k1_g9_x_f1k1 69.51% 69.15% 96.97% 77.41% 23.03% 83.10% 69.25% 2.72% 2.58% 4.75% 13.73% 1144 4 tc_f16k4f2k2f1k1_g9_x_f1k1 69.62% 68.42% 96.45% 82.27% 23.08% 81.68% 69.08% 2.68% 2.52% 4.55% 13.54% 1148 4 ta_f16k4f2k2f1k1_g9_x_f1k1 69.21% 68.84% 97.87% 76.22% 22.77% 81.70% 75.23% 2.65% 2.45% 4.35% 13.76% 1152 4 f1k1_x_g9_f1k1f2k2f4k4f8k8_td 69.62% 67.87% 97.93% 88.79% 23.73% 86.99% 78.63% 2.55% 2.35% 4.15% 12.71% 1205 1 f1k1_x_g9_f1k1f4k2f16k4f64k8_td 69.56% 67.48% 97.42% 86.68% 24.48% 86.35% 78.06% 2.45% 2.25% 3.95% 12.52% 1209 1 f1k1_x_g9_f1k1f8k2f64k4f512k8_td 69.35% 67.89% 97.85% 88.21% 24.66% 76.99% 79.56% 2.35% 2.15% 3.75% 12.13% 1213 1 f1k1_x_g9_f1k1f8k2f64k8f512k16_td 69.20% 68.76% 99.11% 89.18% 24.35% 77.62% 79.88% 2.35% 2.05% 3.55% 11.10% 1239 1 f1k1_x_g9_f1k1f4k2f16k4_td 69.25% 67.40% 98.59% 89.01% 24.24% 77.31% 79.16% 2.45% 1.95% 3.35% 9.22% 1221 1 f1k1_x_g1_f1k1f2k2f16k4_td 69.74% 67.04% 98.09% 79.85% 24.89% 77.27% 80.86% 2.55% 2.15% 3.75% 11.37% 1148 3 f1k1_x_g4_f1k1f2k2f16k4_td 70.28% 68.11% 98.30% 79.45% 24.87% 77.95% 80.23% 2.45% 2.05% 3.75% 11.12% 1164 2 f1k1_x_g9_f1k1f2k2f16k4_td 70.73% 67.97% 98.11% 88.76% 25.79% 78.45% 84.01% 2.35% 1.95% 3.65% 11.98% 1225 1 f1k1_x_g9_f1k1f2k2f16k4_tc 70.48% 68.74% 98.16% 89.18% 27.01% 87.21% 81.04% 2.18% 1.85% 2.54% 11.71% 1229 1 f1k1_x_g9_f1k1f2k2f16k4_ta 70.81% 67.31% 98.15% 80.72% 25.60% 85.60% 82.76% 2.25% 1.77% 2.95% 9.85% 1233 1
4.3.2 Drifting Measurements
We discuss the methods to measure drifting. Several metrics from VBench-Long [19] already integrate long-range consistency measurements and may indicate drifting. We propose a more direct metric, namely start-end contrast, to measure drifting from multiple aspects.
@@ 4.4 Ablative Results
As shown in Table 1, we note several discoveries: (1) The inverted anti-drifting sampling method achieves the best results in 5 out of 7 metrics, while other sampling methods achieve at most a single best metric. (2) The inverted anti-drifting sampling achieves the best performance in all drifting metrics. (3) Human evaluations indicate that generating 9 frames per section yields better perception than generating 1 or 4 frames, as evidenced by the higher ELO scores for configurations with g9 compared to their g1 or g4 counterparts. (4) While vanilla sampling achieved the highest dynamic score, this is likely attributable to drifting effects rather than genuine quality. (5) We also observe that differences between specific configuration options within the same sampling approach are relatively small and random, suggesting that the sampling category contributes more to the overall performance difference.

### v3

@@ 4 Drift Prevention
Figure 2: Anti-drifting sampling and training methods. We present sampling approaches to generate frames in different temporal orders. The shadowed squares are the generated frames in each iteration, whereas the white squares are the iteration inputs. We also discuss the method to convert the frame history into a discrete representation.
4.1 Planned Endpoints and Adjusted Sampling Order
One possible explanation of drifting is that the modeling of the chained conditional probability $\mathbb{P}(\bm{X}_{t}|\bm{X}_{t-1})$ fails to approximate $\mathbb{P}(\bm{X}_{t})$ due to imperfect estimations. This perspective indicates that a strict causal system is more vulnerable to drifting than bi-directional models that directly approximate $\mathbb{P}(\bm{X}_{t}|\bm{X}_{t_{1}},\bm{X}_{t_{2}})$ where $t_{1}<t<t_{2}$.
Endpoint planning
@@ 4.1 Planned Endpoints and Adjusted Sampling Order
The vanilla sampling method shown in Fig. 2-(a) can be modified into Fig. 2-(b), where the first iteration simultaneously generates both beginning and ending sections, while subsequent iterations fill the gaps between these anchors. This simple method can get rid of drifting in specific cases when the video motions are in a relatively small range, or when the motion patterns are repeated or periodic (e.g., dancing, talking, spinning, etc.), or when the motion content follows some texture patterns (e.g., fire flame, water flow, etc.).
Inverted sampling (image-to-video)
@@
A variant by inverting the sampling order in Fig. 2-(b) into Fig. 2-(c) is effective for image-to-video generation. In image-to-video, the first frame is a ground-truth user input, whereas the last frame is a generated endpoint that is not guaranteed to perfectly preserve the quality of the user input. All generations in Fig. 2-(c) keep the direction to approximate the high-quality user input, leading to iteratively refined generations.
Multiple endpoints
The endpoint planning can be repeated with different prompts before filling in the gaps, resulting in a planned sequence of generation (Fig. 2-(d)). Note that though the drifting might still happen in the endpoint-wise, in certain cases, when the prompted sections are distant enough, the error accumulation becomes almost negligible. This method is more flexible than single endpoint planning and supports more dynamic motions and more complicated storytelling.
@@
These sampling methods require modifications to RoPE to support non-consecutive phases (time indices of frames). This is achieved by skipping the blank phases (indices) in the time dimension.
4.2 History Discretization
Another potential cause of drifting is the difference between training and inference distributions over the history frames. This indicates that drifting can be mitigated if the history representation is not sensitive enough to distinguish between the training and inference frames. Discrete integer tokens are well-suited to reduce the mode gap between training and inference distributions. This perspective is supported by empirical evidence that discrete autoregressive systems (e.g., LLMs) often demonstrate less obvious drifting than continuous autoregressive systems for visual content diffusion.
We discretize the history as in Fig. 2-(e). Consider a dataset $\Phi\in\mathbb{R}^{(B\times T\times H\times W)\times C}$ of precomputed latent videos, a $K$-Mean over all latent pixels will yield a codebook $\Omega\in\mathbb{R}^{K\times C}$ where $K\in\mathbb{Z}^{+}$ is an adjustable number. Any latent frame $\bm{F}\in\mathbb{R}^{T\times H\times W\times C}$ can be quantized by $Q(\cdot)$ with
[…]
where $p$ is pixel position, and $Q(\bm{F})\in[0,K-1]^{T\times H\times W}$ is a matrix of indices over the codebook $\Omega$. The matrix of indices can be converted back to latent videos by $\Omega_{Q(\bm{F})}\in\mathbb{R}^{T\times H\times W\times C}$. We replace all history frames $\bm{F}$ with $\Omega_{Q(\bm{F})}$ during training.
Intuitively, when $K=1$, the history becomes meaningless as one single color, and the drifting is eliminated at the cost of giving up memory (errors do not accumulate but sections become unrelated); when $K\to\infty$, the effect is equivalent to no discretization, and the drifting remains. We show with ablation study that a suitable $K$ can minimize the error propagation while simultaneously producing plausible consistency between sections.
@@ 5.1 Ablative Naming
Kernel (simplified): For simplicity, since kernels that are multiples of $(1,2,2)$ are commonly used, we use abbreviated notation such as k1 that only denotes the temporal dimension. Specifically, k1 represents k1h2w2 (the kernel $(1,2,2)$), k2 represents k2h4w4 (the kernel $(2,4,4)$), etc.
@@
Packing: The notation f16k4f2k2f1k1 shows a way to encode $19$ contiguous frames: the first $16$ frames are encoded by the kernel k4 (the kernel $(4,8,8)$), the next $2$ frames are encoded by the kernel k2 (the kernel $(2,4,4)$), and the last $1$ frame is encoded by the kernel k1 (the kernel $(1,2,2)$).
@@
History Discretization: The notation +d means converting history into discrete space.
@@
With the above naming convention, we can represent all ablative structures in a compact form. Note that this naming also implies the sampling approach as discussed in Section 4.1:
@@ 5.2 Base Model and Implementation Details
We implement FramePack with Wan and HunyuanVideo. We implement both the text-to-video and image-to-video structures, though both are naturally supported by next-frame-section prediction models and do not need architecture modifications. We report results with HunyuanVideo in the main paper (see also supplementary for Wan results). We conduct all experiments using H100 GPU clusters with training details in the supplementary. Note that FramePack achieves a batch size of about 64 on a single 8×A100-80G node with the 13B HunyuanVideo model at 480p resolution LoRA training with window size 2 or 3 (or batch size 32 of window size 4 or 5), making FramePack suitable for personal or laboratory-scale training and experimentation. We follow the guidelines of LTXVideo [17]’s dataset collection pipeline to gather data at multiple resolutions and quality levels (see also supplementary for more details).
5.3 Quantitative Evaluation
@@
Metrics
Multiple metrics for video evaluations are consistent with common benchmarks, e.g., VBench [24], VBench2 [75], etc. Clarity: The MUSIQ [27] image quality predictor trained on SPAQ [14]. This metric measures artifacts like noise and blurring. Aesthetic: The LAION aesthetic predictor [40]. This metric measures the aesthetic values perceived by a CLIP-based estimator. Motion: The video frame interpolation model [23] modified by VBench to measure the smoothness of motion. Dynamic: The RAFT [45] modified by VBench to estimate the degree of dynamics. Note that the “dynamic” metric and “motion” metric represent a trade-off, e.g., a still image may rank high on motion smoothness but will be penalized by low dynamic degrees. Semantic: The video-text score computed by ViCLIP [53]. This metric measures the overall semantic consistency between the generated video and the prompt. Anatomy: The ViT [13] pretrained by VBench for identifying the per-frame presence of hands, faces, bodies, etc. Identity: The facial feature similarity using ArcFace [11] with face detection by RetinaFace [12].
Drifting measurement
@@
As shown in Table 1, we note several discoveries. (1) The inverted anti-drifting sampling method achieves the best results in 4 out of 7 metrics, and achieves the best performance in all drifting metrics. (2) However, the inverted anti-drifting sampling has a relatively small dynamic range. (3) While vanilla sampling achieved the highest dynamic score, this is likely attributable to drifting effects rather than genuine quality, evidenced by relatively low ELO scores. (4) The vanilla sampling with discrete history achieves highly competitive human scores while having a much larger dynamic range. (5) We also observe that differences between specific configuration options within the same sampling approach are relatively small and random, suggesting that the overall architecture contributes more to the general difference.
History discretization parameter
The history discretization is influenced by the parameter $K$ with higher $K$ giving stronger anti-drifting effects but also more challenging learning for smooth transitions between sections. In our tests, $K=128$ gives strong drift reduction with relatively minimal training difficulties. We provide more detailed ablations for $K$ in the supplementary materials.
Additional evaluations
We test feature-similarity-based packing with video game (world model) benchmarks in the supplementary material. See also supplementary material for Wan results and more details of decomposed metrics.
Table 1: Ablation study. We evaluate different FramePack configurations across multiple global metrics, drifting metrics, and human assessments. The table is divided into 4 groups based on sampling approach: vanilla sampling, anti-drifting sampling, and inverted anti-drifting sampling, and vanilla sampling with discrete history. The tests are conducted with HunyuanVideo as base. Bests in bold. ELO differences within $\pm 16$ are considered ties. Variant Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank* $\downarrow$ td_f8k8f4k4f2k2f1k1_g9 67.33% 65.94% 94.53% 92.91% 20.06% 66.97% 71.72% 3.25% 3.45% 7.45% 16.56% 1090 5 td_f64k8f16k4f4k2f1k1_g9 67.71% 66.71% 95.09% 93.91% 21.39% 66.56% 71.50% 3.22% 3.38% 7.25% 16.43% 1070 5 td_f512k8f64k4f8k2f1k1_g9 67.74% 66.44% 94.27% 92.91% 21.90% 67.73% 71.35% 3.18% 3.32% 7.05% 15.95% 1085 5 td_f512k16f64k8f8k2f1k1_g9 66.13% 64.52% 94.77% 94.18% 19.21% 65.72% 71.51% 3.25% 3.45% 7.85% 17.62% 1068 5 td_f16k4f4k2f1k1_g9 67.37% 65.05% 95.97% 91.66% 19.15% 65.38% 71.73% 3.22% 3.48% 6.65% 17.36% 1072 5 td_f16k4f2k2f1k1_g1 65.81% 64.39% 94.91% 94.92% 19.20% 64.16% 69.70% 3.43% 3.77% 9.81% 20.89% 1030 6 td_f16k4f2k2f1k1_g4 66.57% 64.99% 94.00% 82.85% 19.40% 65.97% 69.06% 3.36% 3.68% 8.55% 19.09% 1050 6 td_f16k4f2k2f1k1_g9 67.15% 65.29% 94.08% 92.97% 20.73% 66.46% 71.08% 3.18% 3.42% 7.45% 18.05% 1074 5 tc_f16k4f2k2f1k1_g9 67.62% 65.07% 94.02% 90.33% 21.71% 71.70% 71.01% 3.15% 3.25% 7.25% 17.21% 1088 5 ta_f16k4f2k2f1k1_g9 67.02% 66.44% 94.11% 91.09% 21.48% 71.92% 72.18% 3.12% 3.18% 7.05% 17.66% 1092 5 td_f8k8f4k4f2k2f1k1_g9_x_f1k1 68.46% 66.95% 96.46% 75.55% 22.88% 85.10% 75.84% 2.95% 2.85% 5.95% 15.87% 1135 3 td_f64k8f16k4f4k2f1k1_g9_x_f1k1 68.21% 67.75% 95.74% 85.97% 22.79% 81.09% 76.29% 2.92% 2.98% 5.95% 15.99% 1140 3 td_f512k8f64k4f8k2f1k1_g9_x_f1k1 68.62% 67.72% 95.05% 76.50% 22.44% 82.01% 76.65% 2.88% 2.92% 5.75% 15.94% 1118 4 td_f512k16f64k8f8k2f1k1_g9_x_f1k1 68.35% 67.89% 97.73% 76.48% 22.75% 78.40% 76.16% 2.85% 2.85% 5.55% 14.04% 1115 4 td_f16k4f4k2f1k1_g9_x_f1k1 68.42% 67.59% 96.74% 74.37% 23.69% 79.14% 77.37% 2.82% 2.78% 5.35% 14.90% 1138 3 td_f16k4f2k2f1k1_g1_x_f1k1 65.32% 67.97% 95.26% 81.69% 19.97% 74.57% 77.93% 2.78% 2.72% 4.95% 14.80% 1080 5 td_f16k4f2k2f1k1_g4_x_f1k1 69.92% 67.49% 97.84% 71.90% 21.12% 74.84% 77.53% 2.75% 2.65% 4.95% 14.98% 1100 4 td_f16k4f2k2f1k1_g9_x_f1k1 69.51% 69.15% 96.97% 77.41% 23.03% 83.10% 69.25% 2.72% 2.58% 4.75% 13.73% 1142 3 tc_f16k4f2k2f1k1_g9_x_f1k1 69.62% 68.42% 96.45% 82.27% 23.08% 81.68% 69.08% 2.68% 2.52% 4.55% 13.54% 1145 3 ta_f16k4f2k2f1k1_g9_x_f1k1 69.21% 68.84% 97.87% 76.22% 22.77% 81.70% 75.23% 2.65% 2.45% 4.35% 13.76% 1150 3 f1k1_x_g9_f1k1f2k2f4k4f8k8_td 69.62% 67.87% 97.93% 88.79% 23.73% 86.99% 78.63% 2.55% 2.35% 4.15% 12.71% 1210 1 f1k1_x_g9_f1k1f4k2f16k4f64k8_td 69.56% 67.48% 97.42% 86.68% 24.48% 86.35% 78.06% 2.45% 2.25% 3.95% 12.52% 1215 1 f1k1_x_g9_f1k1f8k2f64k4f512k8_td 69.35% 67.89% 97.85% 88.21% 24.66% 76.99% 79.56% 2.35% 2.15% 3.75% 12.13% 1220 1 f1k1_x_g9_f1k1f8k2f64k8f512k16_td 69.20% 68.76% 99.11% 89.18% 24.35% 77.62% 79.88% 2.35% 2.05% 3.55% 11.10% 1235 1 f1k1_x_g9_f1k1f4k2f16k4_td 69.25% 67.40% 98.59% 89.01% 24.24% 77.31% 79.16% 2.45% 1.95% 3.35% 9.22% 1225 1 f1k1_x_g1_f1k1f2k2f16k4_td 69.74% 67.04% 98.09% 79.85% 24.89% 77.27% 80.86% 2.55% 2.15% 3.75% 11.37% 1150 3 f1k1_x_g4_f1k1f2k2f16k4_td 70.28% 68.11% 98.30% 79.45% 24.87% 77.95% 80.23% 2.45% 2.05% 3.75% 11.12% 1175 2 f1k1_x_g9_f1k1f2k2f16k4_td 70.73% 67.97% 98.11% 88.76% 25.79% 78.45% 84.01% 2.35% 1.95% 3.65% 11.98% 1228 1 f1k1_x_g9_f1k1f2k2f16k4_tc 70.48% 68.74% 98.16% 89.18% 27.01% 87.21% 81.04% 2.18% 1.85% 2.54% 11.71% 1230 1 f1k1_x_g9_f1k1f2k2f16k4_ta 70.81% 67.31% 98.15% 80.72% 25.60% 85.60% 82.76% 2.25% 1.77% 2.95% 9.85% 1232 1 td_f8k8f4k4f2k2f1k1_g9+D 69.12% 66.53% 96.91% 93.07% 24.79% 82.58% 73.34% 2.43% 2.74% 4.45% 14.37% 1225 1 td_f64k8f16k4f4k2f1k1_g9+D 69.35% 65.99% 97.21% 92.83% 22.43% 81.88% 72.90% 2.22% 2.44% 4.49% 13.74% 1170 2 td_f512k8f64k4f8k2f1k1_g9+D 69.26% 67.24% 96.28% 91.01% 24.63% 82.58% 71.07% 2.26% 2.01% 5.46% 13.00% 1169 2 td_f512k16f64k8f8k2f1k1_g9+D 69.93% 67.26% 97.60% 93.61% 23.87% 81.27% 73.39% 2.93% 2.92% 5.11% 15.70% 1139 3 td_f16k4f4k2f1k1_g9+D 69.49% 67.25% 96.95% 92.39% 23.40% 88.67% 72.08% 2.30% 2.63% 6.74% 13.51% 1223 1 td_f16k4f2k2f1k1_g1+D 68.16% 65.47% 95.94% 93.15% 23.97% 75.24% 71.78% 2.66% 3.70% 6.40% 16.83% 1145 3 td_f16k4f2k2f1k1_g4+D 68.16% 65.80% 95.26% 92.97% 22.97% 69.87% 71.91% 2.70% 3.73% 5.69% 17.06% 1142 3 td_f16k4f2k2f1k1_g9+D 69.78% 67.22% 96.91% 91.39% 24.40% 77.62% 74.95% 2.30% 2.49% 4.12% 14.11% 1222 1 tc_f16k4f2k2f1k1_g9+D 69.42% 66.24% 96.01% 92.53% 23.06% 79.15% 72.76% 2.87% 2.43% 5.51% 13.67% 1218 1 ta_f16k4f2k2f1k1_g9+D 68.51% 67.77% 96.90% 93.92% 23.24% 79.90% 70.01% 2.27% 2.41% 4.31% 15.85% 1216 1 * Based on the ELO scores and tie rules, the ranks are divided into: 1030-1050 (Rank 6), 1068-1092 (Rank 5), 1100-1118 (Rank 4), 1135-1150 (Rank 3), 1169-1175 (Rank 2), and 1210-1235 (Rank 1).